# A class of forward-backward regularizations of the Perona-Malik equation with variable exponent

### v2

@@ 2.1 Variable exponent spaces
There are many examples of John domains. Every ball is clearly a 1-John domain, and every bounded domain satisfying the uniform interior cone condition is a John domain. In particular, any bounded Lipschitz domain is a John domain, where a domain with $C^{0,1}$-boundary is referred to as a Lipschitz domain; see also [Diening2011, Martio1979383].
For later use, we state the following important properties of the spaces $L^{p(\cdot)}(\Omega)$ and $W^{1,p(\cdot)}(\Omega)$. Detailed discussions can be found in [Diening2011, Kovavcik1991592, Mihuailescu20072929].
Lemma 1
[Diening2011] Let $p\in\mathcal{P}(\Omega)$ with $p^{-}<\infty$. If $\varrho_{L^{p(\cdot)}(\Omega)}(u)>0$ or $p^{+}<\infty$, then
@@
Lemma 2
[Diening2011] Let $p\in\mathcal{P}(\Omega)$. Then $L^{p(\cdot)}(\Omega)$ and $W^{k,p(\cdot)}(\Omega)$ are Banach spaces, separable if $p$ is bounded, and reflexive and uniformly convex whenever $1<p^{-}\leq p^{+}<\infty$.
Lemma 3 (Hölder’s inequality)
[Diening2011] Let $p,q,s\in\mathcal{P}(\Omega)$ be such that $\frac{1}{s(y)}=\frac{1}{p(y)}+\frac{1}{q(y)}$ for a.e. $y\in\Omega$. Then, for all $u\in L^{p(\cdot)}(\Omega)$ and $v\in L^{q(\cdot)}(\Omega)$,
@@
Lemma 4
[Diening2011] Let $\Omega$ be a Lipschitz domain, and let $p\in\mathcal{P}^{\log}(\Omega)$ satisfy $1\leq p^{-}\leq p^{+}<\infty$. Then $C^{\infty}(\bar{\Omega})$ is dense in $W^{k,p(\cdot)}(\Omega)$, $k\in\mathbb{N}$.
Lemma 5 (Poincaré inequality)
[Diening2011] Let $p\in\mathcal{P}^{\log}(\Omega)$. If $\Omega$ is a bounded $\alpha$-John domain, then
@@
By the weak $\ast$ convergence of $\{f^{m}\}_{m=1}^{\infty}$ (see [Zeidler1990]), there exists $N_{i}\in\mathbb{N}$ such that, for all $m>N_{i}$, it holds that
@@
Fix $N=\max_{1\leq i\leq n}N_{i}$. For any $\xi\in\varPsi$, there exists $j\in\{1,2,\cdots,n\}$ such that $\left\|\xi-\xi_{j}\right\|_{L^{1}(0,T;L^{p^{\prime}(\cdot)}(\Omega))}<r$. Consequently, for all $m>N$, it follows from Lemma 3 that
@@
This completes the proof.
Adapting the argument of Theorem 3.20 in [Dacorogna2008], the following sufficient condition for weak lower semicontinuity on $W^{1,p(\cdot)}(\Omega)$ can be derived. The proof is omitted for brevity.
Lemma 7
@@ 2.2 Young measures
Lemma 8
[ball2005version] Let $\Omega\subset\mathbb{R}^{N}$ be Lebesgue measurable, let $K\subset\mathbb{R}^{N}$ be closed, and let $u_{j}:\Omega\rightarrow\mathbb{R}^{N},j\in\mathbb{N}$, be a sequence of Lebesgue measurable functions satisfying $u_{j}\rightarrow K$ in measure as $j\rightarrow\infty$, i.e., given any open neighbourhood $U$ of $K$ in $\mathbb{R}^{N}$
@@
$\forall R>0:~\lim_{L\rightarrow\infty}\sup_{k\in\mathbb{N}}\left|\{x\in\Omega\cap B_{R}:|u_{j_{k}}(x)|\geq L\}\right|=0,$ (14)
@@
$\left\|\nu_{x}\right\|_{\mathcal{M}(\mathbb{R}^{N})}=1\quad\text{ for a.e. }x\in\Omega,$ (15)
@@
Definition 3
The map $\nu:\Omega\rightarrow\mathcal{M}(\mathbb{R}^{N})$ in Lemma 8 is called the Young measure in $\mathbb{R}^{N}$ generated by the sequence $\{u_{j_{k}}\}_{k=1}^{\infty}$.
Hungerbühler subsequently established the following consequence, while refined versions of Lemma 8 can be found in [Kinderlehrer1994059].
Lemma 9
[Hungerbuhler1997] Let $\Omega$, $u_{j}$, and $\nu_{x}$ be as in Lemma 8. Then (14), (15) and (16) are equivalent.
Remark 1
In [ball2005version], it is also shown that, under hypothesis (14), for any measurable set $D\subset\Omega$,
@@
for every Carathéodory function $f:D\times\mathbb{R}^{N}\rightarrow\mathbb{R}$ such that $\left\{f(\cdot,u_{j_{k}})\right\}_{k=1}^{\infty}$ is sequentially weakly relatively compact in $L^{1}(D)$. This statement is thus also equivalent to (14), (15), and (16).
2.3 Definition of Young measure solutions
We define a Young measure solution to problem (1)–(3). For clarity, let $\rho\in W^{1,p(\cdot)}(\Omega)$ be fixed but arbitrary, and define
@@ 2.3 Definition of Young measure solutions
as the affine subspace of $W^{1,p(\cdot)}(\Omega)$. Recall that the variable exponent $p$ satisfies condition (4), and denote by
@@
the heat flux in equation (1), where $\vec{q}(x,\xi)=\nabla_{2}\varphi(x,\xi)$ and $\nabla_{2}$ denotes the derivative with respect to the second variable. The potential $\varphi:\Omega\times\mathbb{R}^{N}\rightarrow\mathbb{R}$ is a Carathéodory function, nonconvex with respect to $\xi$ for sufficiently small $\delta$. It is noted that $\varphi(x,\cdot)\in C^{2}(\mathbb{R}^{N})$ for all $x\in\Omega$, and that $\vec{q}$ and $\varphi$ satisfy the following growth conditions:
$\left|\vec{q}(x,\xi)\right|\leq\lambda_{2}\left|\xi\right|^{p(x)-1}+1,\quad\max\left\{\lambda_{1}\left|\xi\right|^{p(x)}-1,0\right\}\leq\varphi(x,\xi)\leq\lambda_{2}\left|\xi\right|^{p(x)}+1,$ (17)
@@
$\left(\vec{q}(x,\xi_{1})-\vec{q}(x,\xi_{2})\right)\cdot\left(\xi_{1}-\xi_{2}\right)\geq-K\left|\xi_{1}-\xi_{2}\right|^{2},$ (18)
@@
Definition 4
A pair $(u,\nu)$ is called a Young measure (or measure-valued) solution of problem (1)–(3) if $u\in L^{\infty}(0,T;W_{f}^{1,p(\cdot)}(\Omega))\cap H^{1}(0,T;L^{2}(\Omega))$ and $\nu=(\nu_{x,t})_{(x,t)\in Q_{T}}$ is a parametrized family of probability measures on $\mathbb{R}^{N}$, such that $u(0)=f$ a.e. in $\Omega$, and
$\nabla u(x,t)=\langle\nu_{x,t},\mathrm{id}\rangle=\int_{\mathbb{R}^{N}}\xi d\nu_{x,t}(\xi)\quad\text{a.e. in }Q_{T},$ (19)
$\iint_{Q_{T}}\left(u_{t}\zeta+\langle\nu_{x,t},\vec{q}(x,\cdot)\rangle\cdot\nabla\zeta\right)dxdt=0,$ (20)
@@
Our goal is to construct a Young measure solution to problem (1)–(3). Following the approach in [Oliver20031380, THANH20141403], we first consider the homogeneous Neumann IBVP for the Sobolev regularization of equation (1):
$\displaystyle u^{\varepsilon}_{t}=\operatorname{div}\vec{q}(x,\nabla u^{\varepsilon})+\varepsilon\Delta u^{\varepsilon}_{t},$ $(x,t)\in Q_{T}$, (21)
$\displaystyle\left(\vec{q}\left(x,\nabla u^{\varepsilon}\right)+\varepsilon\nabla u_{t}^{\varepsilon}\right)\cdot\vec{n}=0,$ $(x,t)\in\partial\Omega\times(0,T)$, (22)
$\displaystyle u^{\varepsilon}(x,0)=f,$ $x\in\Omega$, (23)
where $0<\varepsilon<1$. We refer to the problem (21)–(23) as the regularized problem corresponding to the original problem (1)–(3).
3 Weak solutions of the regularized problem (21)–(23)
In this section, we study the well-posedness of problem (21)–(23) and establish the following theorem by employing Rothe’s method and variational principles.
Theorem 2
Let $f\in W^{1,p(\cdot)}(\Omega)$. Then the regularized problem (21)–(23) admits a unique weak solution $u^{\varepsilon}\in L^{\infty}(0,T;W_{f}^{1,p(\cdot)}(\Omega))\cap H^{1}(0,T;H^{1}(\Omega))$, satisfying $u^{\varepsilon}(0)=f$ a.e. in $\Omega$, and
$\iint_{Q_{T}}\left(\frac{\partial u^{\varepsilon}}{\partial t}\zeta+\vec{q}(x,\nabla u^{\varepsilon})\cdot\nabla\zeta+\varepsilon\nabla\frac{\partial u^{\varepsilon}}{\partial t}\cdot\nabla\zeta\right)dxdt=0,$ (24)
@@ 3 Weak solutions of the regularized problem (21)–(23)
$\int_{\Omega}\varphi(x,\nabla u^{\varepsilon}(t))dx+\left\|\frac{\partial u^{\varepsilon}}{\partial\tau}\right\|^{2}_{L^{2}(Q_{t})}+\left\|\nabla\sqrt{\varepsilon}\frac{\partial u^{\varepsilon}}{\partial\tau}\right\|^{2}_{L^{2}(Q_{t})}\leq\int_{\Omega}\varphi(x,\nabla f)dx,$ (25)
@@
Remark 2
By substituting the constant test function $\zeta=1_{\Omega}$ into the weak formulation (25), we obtain
@@
Therefore, it is natural to require that the solution to problem (21)–(23) satisfies $u^{\varepsilon}(\cdot,t)\in W_{f}^{1,p(\cdot)}(\Omega)$ for a.e. $t\in[0,T]$.
@@
where $u_{m}^{0}=f$, $u_{m}^{j-1}\in W_{f}^{1,p(\cdot)}(\Omega)$, $j=1,2,\cdots,m$. The energy functional above is lower bounded due to the growth condition (17). It is readily seen that for every $v\in W^{1,p(\cdot)}(\Omega)$,
@@
It is clear that $u\mapsto J^{m,j}(u)-\int_{\Omega}\varphi_{\varepsilon}(x,\nabla u)dx$ is weakly lower semicontinuous on $W^{1,p(\cdot)}(\Omega)$. Additionally, by (18),
$\left(\nabla_{2}\varphi_{\varepsilon}(x,\xi_{1})-\nabla_{2}\varphi_{\varepsilon}(x,\xi_{2})\right)\cdot\left(\xi_{1}-\xi_{2}\right)\geq\left(-K+\varepsilon\frac{m}{T}\right)\left|\xi_{1}-\xi_{2}\right|^{2},\quad x\in\Omega,~\xi_{1},\xi_{2}\in\mathbb{R}^{N},$ (26)
which shows that the mapping $\xi\mapsto\varphi_{\varepsilon}(x,\xi)$ is convex for every $x\in\Omega$ whenever $m\geq\frac{KT}{\varepsilon}$. Lemma 7 then yields the weak lower semicontinuity of $u\mapsto\int_{\Omega}\varphi_{\varepsilon}(x,\nabla u)dx$ on $W^{1,p(\cdot)}(\Omega)$. Let $\{u_{m}^{j,k}\}_{k=1}^{\infty}\subset W_{f}^{1,p(\cdot)}(\Omega)$ be a minimizing sequence of $J^{m,j}$. Since $J^{m,j}$ is coercive on $W_{f}^{1,p(\cdot)}(\Omega)$ by Lemma 5, we may assume that $\{u_{m}^{j,k}\}_{k=1}^{\infty}$ is bounded in $W^{1,p(\cdot)}(\Omega)$. Applying Lemma 2, there exists $u_{m}^{j}\in W_{f}^{1,p(\cdot)}(\Omega)$ and a subsequence of $\{u_{m}^{j,k}\}_{k=1}^{\infty}$, still denoted by itself, such that $u_{m}^{j,k}\rightharpoonup u_{m}^{j}$ weakly in $W^{1,p(\cdot)}(\Omega)$. Then, we have
@@
Consequently, $u_{m}^{j}\in W_{f}^{1,p(\cdot)}(\Omega)$ is a (not necessarily unique) minimizer of $J^{m,j}$. This finishes the proof.
@@
$\int_{\Omega}\left(\frac{m}{T}\left(u_{m}^{j}-u_{m}^{j-1}\right)\phi+\vec{q}(x,\nabla u_{m}^{j})\cdot\nabla\phi+\varepsilon\frac{m}{T}\left(\nabla u_{m}^{j}-\nabla u_{m}^{j-1}\right)\cdot\nabla\phi\right)dx=0.$ (27)
From (17) and the facts that $\nabla u_{m}^{j}\in L^{p(\cdot)}(\Omega)$ and $\vec{q}(x,\nabla u_{m}^{j})\in L^{p^{\prime}(\cdot)}(\Omega)$, together with Lemma 4, it follows that $\phi\in W^{1,p(\cdot)}(\Omega)$ is admissible as a test function in the the Euler-Lagrange system (27) for $u_{m}^{j}$. We now derive an a priori estimate for the discrete energy.
Lemma 11 (Discrete energy inequality)
@@
$\int_{\Omega}\varphi(x,\nabla u_{m}^{j})dx+\sum_{i=1}^{j}\int_{\Omega}\frac{m}{T}\left|u_{m}^{i}-u_{m}^{i-1}\right|^{2}dx+\sum_{i=1}^{j}\int_{\Omega}(\varepsilon-\gamma)\frac{m}{T}\left|\nabla u_{m}^{i}-\nabla u_{m}^{i-1}\right|^{2}dx\leq\int_{\Omega}\varphi(x,\nabla f)dx,$ (28)
@@
Proof 3
Similar to the treatment in (26), it follows from (18) that, for $m\geq\frac{KT}{2\gamma}$, the mapping $\xi\mapsto\varphi(x,\xi)+\gamma\frac{m}{T}|\xi-\nabla u_{m}^{j-1}|^{2}$, $j=1,2,\cdots,m$, is convex; i.e., for all $x\in\Omega$ and $\xi_{1},\xi_{2}\in\mathbb{R}^{N}$, we have
@@
$\leq\left(\vec{q}(x,\xi_{1})+2\gamma\frac{m}{T}\left(\xi_{1}-\nabla u_{m}^{j-1}(x)\right)\right)\cdot\left(\xi_{1}-\xi_{2}\right).$ (29)
Set $m_{0}=\max\left\{\frac{KT}{2\gamma},\frac{KT}{\varepsilon}\right\}$. Then, by Lemma 10, $J^{m,j}$ admits a minimizer $u_{m}^{j}\in W_{f}^{1,p(\cdot)}(\Omega)$, satisfying (27) for every $m\geq m_{0}$ and $j=1,2,\cdots,m$. From (29), for $i=1,2,\cdots,j$, we derive
@@
$-(\varepsilon-\gamma)\frac{m}{T}\left|\nabla u_{m}^{i}-\nabla u_{m}^{i-1}\right|^{2}\bigg]dx.$ (30)
Before proceeding with the estimate, we consider (27) with $\phi=u_{m}^{i}-u_{m}^{i-1}$, which yields
@@
We insert the above equality into (30) to obtain
@@
This completes the proof.
Corollary 1
@@
$\sup_{1\leq j\leq m}\left\|u_{m}^{j}\right\|_{W^{1,p(\cdot)}(\Omega)}+\sum_{j=1}^{m}\frac{m}{T}\left\|u_{m}^{j}-u_{m}^{j-1}\right\|_{L^{2}(\Omega)}^{2}+\sum_{j=1}^{m}\left(\varepsilon-\gamma\right)\frac{m}{T}\left\|\nabla u_{m}^{j}-\nabla u_{m}^{j-1}\right\|_{L^{2}(\Omega)}^{2}\leq M_{1},$ (31)
@@
The assertion follows directly from the growth condition (17), Lemma 5 and 11, and the assumption that $f\in W^{1,p(\cdot)}(\Omega)$.
Now we construct approximate solutions to the regularized problem (21)–(23). Let $m\geq m_{0}$. For $j=1,2,\cdots,m$, let $\chi_{m}^{j}(t)$ denote the characteristic function of the interval $(j-1)\frac{T}{m},j\frac{T}{m})$, and define $\lambda_{m}^{j}(t)$ as

### v3

@@ 2.1. Variable exponent spaces
There are many examples of John domains. Every ball is clearly a 1-John domain, and every bounded domain satisfying the uniform interior cone condition is a John domain. In particular, any bounded Lipschitz domain is a John domain, where a domain with $C^{0,1}$-boundary is referred to as a Lipschitz domain; see also [18, 38].
For later use, we state the following important properties of the spaces $L^{p(\cdot)}(\Omega)$ and $W^{1,p(\cdot)}(\Omega)$. Detailed discussions can be found in [18, 35, 40].
Lemma 2.3.
[18] Let $p\in\mathcal{P}(\Omega)$ with $p^{-}<\infty$. If $\varrho_{L^{p(\cdot)}(\Omega)}(u)>0$ or $p^{+}<\infty$, then
@@
Lemma 2.4.
[18] Let $p\in\mathcal{P}(\Omega)$. Then $L^{p(\cdot)}(\Omega)$ and $W^{k,p(\cdot)}(\Omega)$ are Banach spaces, separable if $p$ is bounded, and reflexive and uniformly convex whenever $1<p^{-}\leq p^{+}<\infty$.
Lemma 2.5 (Hölder’s inequality).
[18] Let $p,q,s\in\mathcal{P}(\Omega)$ be such that $\frac{1}{s(y)}=\frac{1}{p(y)}+\frac{1}{q(y)}$ for a.e. $y\in\Omega$. Then, for all $u\in L^{p(\cdot)}(\Omega)$ and $v\in L^{q(\cdot)}(\Omega)$,
@@
Lemma 2.6.
[18] Let $\Omega$ be a Lipschitz domain, and let $p\in\mathcal{P}^{\log}(\Omega)$ satisfy $1\leq p^{-}\leq p^{+}<\infty$. Then $C^{\infty}(\bar{\Omega})$ is dense in $W^{k,p(\cdot)}(\Omega)$, $k\in\mathbb{N}$.
Lemma 2.7 (Poincaré inequality).
[18] Let $p\in\mathcal{P}^{\log}(\Omega)$. If $\Omega$ is a bounded $\alpha$-John domain, then
@@
By the weak $\ast$ convergence of $\{f^{m}\}_{m=1}^{\infty}$ (see [56]), there exists $N_{i}\in\mathbb{N}$ such that, for all $m>N_{i}$, it holds that
@@
Fix $N=\max_{1\leq i\leq n}N_{i}$. For any $\xi\in\varPsi$, there exists $j\in\{1,2,\cdots,n\}$ such that $\left\|\xi-\xi_{j}\right\|_{L^{1}(0,T;L^{p^{\prime}(\cdot)}(\Omega))}<r$. Consequently, for all $m>N$, it follows from Lemma 2.5 that
@@
This completes the proof. ∎
Adapting the argument of Theorem 3.20 in [16], the following sufficient condition for weak lower semicontinuity on $W^{1,p(\cdot)}(\Omega)$ can be derived. The proof is omitted for brevity.
Lemma 2.9.
@@ 2.2. Young measures
Lemma 2.10.
[1] Let $\Omega\subset\mathbb{R}^{N}$ be Lebesgue measurable, let $K\subset\mathbb{R}^{N}$ be closed, and let $u_{j}:\Omega\rightarrow\mathbb{R}^{N},j\in\mathbb{N}$, be a sequence of Lebesgue measurable functions satisfying $u_{j}\rightarrow K$ in measure as $j\rightarrow\infty$, i.e., given any open neighbourhood $U$ of $K$ in $\mathbb{R}^{N}$
@@
$\forall R>0:~\lim_{L\rightarrow\infty}\sup_{k\in\mathbb{N}}\left|\{x\in\Omega\cap B_{R}:|u_{j_{k}}(x)|\geq L\}\right|=0,$ (2.1)
@@
$\left\|\nu_{x}\right\|_{\mathcal{M}(\mathbb{R}^{N})}=1\quad\text{ for a.e. }x\in\Omega,$ (2.2)
@@
Definition 2.11.
The map $\nu:\Omega\rightarrow\mathcal{M}(\mathbb{R}^{N})$ in Lemma 2.10 is called the Young measure in $\mathbb{R}^{N}$ generated by the sequence $\{u_{j_{k}}\}_{k=1}^{\infty}$.
Hungerbühler subsequently established the following consequence, while refined versions of Lemma 2.10 can be found in [34].
Lemma 2.12.
[30] Let $\Omega$, $u_{j}$, and $\nu_{x}$ be as in Lemma 2.10. Then (2.1), (2.2) and (2.3) are equivalent.
Remark 2.13.
In [1], it is also shown that, under hypothesis (2.1), for any measurable set $D\subset\Omega$,
@@
for every Carathéodory function $f:D\times\mathbb{R}^{N}\rightarrow\mathbb{R}$ such that $\left\{f(\cdot,u_{j_{k}})\right\}_{k=1}^{\infty}$ is sequentially weakly relatively compact in $L^{1}(D)$. This statement is thus also equivalent to (2.1), (2.2), and (2.3).
2.3. Definition of Young measure solutions
We define a Young measure solution to problem (1.1)–(1.3). For clarity, let $\rho\in W^{1,p(\cdot)}(\Omega)$ be fixed but arbitrary, and define
@@ 2.3. Definition of Young measure solutions
as the affine subspace of $W^{1,p(\cdot)}(\Omega)$. Recall that the variable exponent $p$ satisfies condition (1.4), and denote by
@@
the heat flux in equation (1.1), where $\vec{q}(x,\xi)=\nabla_{2}\varphi(x,\xi)$ and $\nabla_{2}$ denotes the derivative with respect to the second variable. The potential $\varphi:\Omega\times\mathbb{R}^{N}\rightarrow\mathbb{R}$ is a Carathéodory function, nonconvex with respect to $\xi$ for sufficiently small $\delta$. It is noted that $\varphi(x,\cdot)\in C^{2}(\mathbb{R}^{N})$ for all $x\in\Omega$, and that $\vec{q}$ and $\varphi$ satisfy the following growth conditions:
$\left|\vec{q}(x,\xi)\right|\leq\lambda_{2}\left|\xi\right|^{p(x)-1}+1,\quad\max\left\{\lambda_{1}\left|\xi\right|^{p(x)}-1,0\right\}\leq\varphi(x,\xi)\leq\lambda_{2}\left|\xi\right|^{p(x)}+1,$ (2.4)
@@
$\left(\vec{q}(x,\xi_{1})-\vec{q}(x,\xi_{2})\right)\cdot\left(\xi_{1}-\xi_{2}\right)\geq-K\left|\xi_{1}-\xi_{2}\right|^{2},$ (2.5)
@@
Definition 2.14.
A pair $(u,\nu)$ is called a Young measure (or measure-valued) solution of problem (1.1)–(1.3) if $u\in L^{\infty}(0,T;W_{f}^{1,p(\cdot)}(\Omega))\cap H^{1}(0,T;L^{2}(\Omega))$ and $\nu=(\nu_{x,t})_{(x,t)\in Q_{T}}$ is a parametrized family of probability measures on $\mathbb{R}^{N}$, such that $u(0)=f$ a.e. in $\Omega$, and
$\nabla u(x,t)=\langle\nu_{x,t},\mathrm{id}\rangle=\int_{\mathbb{R}^{N}}\xi d\nu_{x,t}(\xi)\quad\text{a.e. in }Q_{T},$ (2.6)
$\iint_{Q_{T}}\left(u_{t}\zeta+\langle\nu_{x,t},\vec{q}(x,\cdot)\rangle\cdot\nabla\zeta\right)dxdt=0,$ (2.7)
@@
Our goal is to construct a Young measure solution to problem (1.1)–(1.3). Following the approach in [44, 51], we first consider the homogeneous Neumann IBVP for the Sobolev regularization of equation (1.1):
$\displaystyle u^{\varepsilon}_{t}=\operatorname{div}\vec{q}(x,\nabla u^{\varepsilon})+\varepsilon\Delta u^{\varepsilon}_{t},$ $(x,t)\in Q_{T}$, (2.8)
$\displaystyle\left(\vec{q}\left(x,\nabla u^{\varepsilon}\right)+\varepsilon\nabla u_{t}^{\varepsilon}\right)\cdot\vec{n}=0,$ $(x,t)\in\partial\Omega\times(0,T)$, (2.9)
$\displaystyle u^{\varepsilon}(x,0)=f,$ $x\in\Omega$, (2.10)
where $0<\varepsilon<1$. We refer to the problem (2.8)–(2.10) as the regularized problem corresponding to the original problem (1.1)–(1.3).
3. Weak solutions of the regularized problem (2.8)–(2.10)
In this section, we study the well-posedness of problem (2.8)–(2.10) and establish the following theorem by employing Rothe’s method and variational principles.
Theorem 3.1.
Let $f\in W^{1,p(\cdot)}(\Omega)$. Then the regularized problem (2.8)–(2.10) admits a unique weak solution $u^{\varepsilon}\in L^{\infty}(0,T;W_{f}^{1,p(\cdot)}(\Omega))\cap H^{1}(0,T;H^{1}(\Omega))$, satisfying $u^{\varepsilon}(0)=f$ a.e. in $\Omega$, and
$\iint_{Q_{T}}\left(\frac{\partial u^{\varepsilon}}{\partial t}\zeta+\vec{q}(x,\nabla u^{\varepsilon})\cdot\nabla\zeta+\varepsilon\nabla\frac{\partial u^{\varepsilon}}{\partial t}\cdot\nabla\zeta\right)dxdt=0,$ (3.1)
@@
$\int_{\Omega}\varphi(x,\nabla u^{\varepsilon}(t))dx+\left\|\frac{\partial u^{\varepsilon}}{\partial\tau}\right\|^{2}_{L^{2}(Q_{t})}+\left\|\nabla\sqrt{\varepsilon}\frac{\partial u^{\varepsilon}}{\partial\tau}\right\|^{2}_{L^{2}(Q_{t})}\leq\int_{\Omega}\varphi(x,\nabla f)dx,$ (3.2)
@@
Remark 3.2.
By substituting the constant test function $\zeta=1_{\Omega}$ into the weak formulation (3.2), we obtain
@@
Therefore, it is natural to require that the solution to problem (2.8)–(2.10) satisfies $u^{\varepsilon}(\cdot,t)\in W_{f}^{1,p(\cdot)}(\Omega)$ for a.e. $t\in[0,T]$.
@@
where $u_{m}^{0}=f$, $u_{m}^{j-1}\in W_{f}^{1,p(\cdot)}(\Omega)$, $j=1,2,\cdots,m$. The energy functional above is lower bounded due to the growth condition (2.4). It is readily seen that for every $v\in W^{1,p(\cdot)}(\Omega)$,
@@
It is clear that $u\mapsto J^{m,j}(u)-\int_{\Omega}\varphi_{\varepsilon}(x,\nabla u)dx$ is weakly lower semicontinuous on $W^{1,p(\cdot)}(\Omega)$. Additionally, by (2.5),
$\left(\nabla_{2}\varphi_{\varepsilon}(x,\xi_{1})-\nabla_{2}\varphi_{\varepsilon}(x,\xi_{2})\right)\cdot\left(\xi_{1}-\xi_{2}\right)\geq\left(-K+\varepsilon\frac{m}{T}\right)\left|\xi_{1}-\xi_{2}\right|^{2},\quad x\in\Omega,~\xi_{1},\xi_{2}\in\mathbb{R}^{N},$ (3.3)
which shows that the mapping $\xi\mapsto\varphi_{\varepsilon}(x,\xi)$ is convex for every $x\in\Omega$ whenever $m\geq\frac{KT}{\varepsilon}$. Lemma 2.9 then yields the weak lower semicontinuity of $u\mapsto\int_{\Omega}\varphi_{\varepsilon}(x,\nabla u)dx$ on $W^{1,p(\cdot)}(\Omega)$. Let $\{u_{m}^{j,k}\}_{k=1}^{\infty}\subset W_{f}^{1,p(\cdot)}(\Omega)$ be a minimizing sequence of $J^{m,j}$. Since $J^{m,j}$ is coercive on $W_{f}^{1,p(\cdot)}(\Omega)$ by Lemma 2.7, we may assume that $\{u_{m}^{j,k}\}_{k=1}^{\infty}$ is bounded in $W^{1,p(\cdot)}(\Omega)$. Applying Lemma 2.4, there exists $u_{m}^{j}\in W_{f}^{1,p(\cdot)}(\Omega)$ and a subsequence of $\{u_{m}^{j,k}\}_{k=1}^{\infty}$, still denoted by itself, such that $u_{m}^{j,k}\rightharpoonup u_{m}^{j}$ weakly in $W^{1,p(\cdot)}(\Omega)$. Then, we have
@@
Consequently, $u_{m}^{j}\in W_{f}^{1,p(\cdot)}(\Omega)$ is a (not necessarily unique) minimizer of $J^{m,j}$. This finishes the proof. ∎
@@
$\int_{\Omega}\left(\frac{m}{T}\left(u_{m}^{j}-u_{m}^{j-1}\right)\phi+\vec{q}(x,\nabla u_{m}^{j})\cdot\nabla\phi+\varepsilon\frac{m}{T}\left(\nabla u_{m}^{j}-\nabla u_{m}^{j-1}\right)\cdot\nabla\phi\right)dx=0.$ (3.4)
From (2.4) and the facts that $\nabla u_{m}^{j}\in L^{p(\cdot)}(\Omega)$ and $\vec{q}(x,\nabla u_{m}^{j})\in L^{p^{\prime}(\cdot)}(\Omega)$, together with Lemma 2.6, it follows that $\phi\in W^{1,p(\cdot)}(\Omega)$ is admissible as a test function in the the Euler-Lagrange system (3.4) for $u_{m}^{j}$. We now derive an a priori estimate for the discrete energy.
Lemma 3.4 (Discrete energy inequality).
@@
$\int_{\Omega}\varphi(x,\nabla u_{m}^{j})dx+\sum_{i=1}^{j}\int_{\Omega}\frac{m}{T}\left|u_{m}^{i}-u_{m}^{i-1}\right|^{2}dx+\sum_{i=1}^{j}\int_{\Omega}(\varepsilon-\gamma)\frac{m}{T}\left|\nabla u_{m}^{i}-\nabla u_{m}^{i-1}\right|^{2}dx\leq\int_{\Omega}\varphi(x,\nabla f)dx,$ (3.5)
@@
Proof.
Similar to the treatment in (3.3), it follows from (2.5) that, for $m\geq\frac{KT}{2\gamma}$, the mapping $\xi\mapsto\varphi(x,\xi)+\gamma\frac{m}{T}|\xi-\nabla u_{m}^{j-1}|^{2}$, $j=1,2,\cdots,m$, is convex; i.e., for all $x\in\Omega$ and $\xi_{1},\xi_{2}\in\mathbb{R}^{N}$, we have
@@
$\leq\left(\vec{q}(x,\xi_{1})+2\gamma\frac{m}{T}\left(\xi_{1}-\nabla u_{m}^{j-1}(x)\right)\right)\cdot\left(\xi_{1}-\xi_{2}\right).$ (3.6)
Set $m_{0}=\max\left\{\frac{KT}{2\gamma},\frac{KT}{\varepsilon}\right\}$. Then, by Lemma 3.3, $J^{m,j}$ admits a minimizer $u_{m}^{j}\in W_{f}^{1,p(\cdot)}(\Omega)$, satisfying (3.4) for every $m\geq m_{0}$ and $j=1,2,\cdots,m$. From (3.6), for $i=1,2,\cdots,j$, we derive
@@
$-(\varepsilon-\gamma)\frac{m}{T}\left|\nabla u_{m}^{i}-\nabla u_{m}^{i-1}\right|^{2}\bigg]dx.$ (3.7)
Before proceeding with the estimate, we consider (3.4) with $\phi=u_{m}^{i}-u_{m}^{i-1}$, which yields
@@
We insert the above equality into (3.7) to obtain
@@
This completes the proof. ∎
Corollary 3.5.
@@
$\sup_{1\leq j\leq m}\left\|u_{m}^{j}\right\|_{W^{1,p(\cdot)}(\Omega)}+\sum_{j=1}^{m}\frac{m}{T}\left\|u_{m}^{j}-u_{m}^{j-1}\right\|_{L^{2}(\Omega)}^{2}+\sum_{j=1}^{m}\left(\varepsilon-\gamma\right)\frac{m}{T}\left\|\nabla u_{m}^{j}-\nabla u_{m}^{j-1}\right\|_{L^{2}(\Omega)}^{2}\leq M_{1},$ (3.8)
@@
The assertion follows directly from the growth condition (2.4), Lemma 2.7 and 3.4, and the assumption that $f\in W^{1,p(\cdot)}(\Omega)$.
Now we construct approximate solutions to the regularized problem (2.8)–(2.10). Let $m\geq m_{0}$. For $j=1,2,\cdots,m$, let $\chi_{m}^{j}(t)$ denote the characteristic function of the interval $(j-1)\frac{T}{m},j\frac{T}{m})$, and define $\lambda_{m}^{j}(t)$ as